Find two consecutive odd integers such that 5 times the first integer is 12 more than 3 times the second.
Answer by checkley71(8403)

LET X & X+2 BE THE 2 ODD INTEGERS.
5X=12+3(X+2)
5X=12+3X+6
5X-3X=18
2X=18
X=18/2
X=9 ANSWER FOR THE SMALLER ODD INTEGER.
9+2=11 FOR THE LARGER ODD INTEGER.
PROOF
5*9=12+3*11
45=12+33
45=45